Operations Report Dec 04th – 08th

MDRS Operations Report 4-8- Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 49.73

Minimum 44.60

Maximum 58.40

Notes on solar: We are adjusting power use to compensate for the weather.

Propane Reading, station tank – FULL %

Propane Reading, director tank— FULL %

Propane Reading, intern tank— 50 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 500 gallons

Static to Loft Pump used – yes

Water in GreenHab – 250 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 550 gallons

Hab toilet tank emptied: no

Perseverance rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Spirit rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Opportunity rover used: No

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Curiosity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Notes on rovers: Percy should be used to short range EVA’s only as the batteries are not at strong as the remaining rovers, which all have new batteries.

ATV’s Used: (Honda, 350.1, 350.2, 300): The tires have been successfully mounted on the rear of the Honda and it is once again operational.

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water

CrewCar used and why, where? Yes, To Grand Junction. Battery does not have enough cranking power in the below freezing mornings.

Luna used and why, where? Yes, to Loa for supplies for the astrobiology project

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Hab is fully operational with 4 people staying there.

Summary of Outpost operations: We have been working on the intern trailer. Painting was completed today, except for the area around the bunks and the bunks themselves. That will need to be done once there are fewer people so the bunk paint can harden before use. New mattress installed on the lower bunk, curtain installed between bathroom and bedroom. Carpet removed in the living area. Tomorrow we will complete the cleanup and the interns and Atila will move in on Thursday morning.

Summary of GreenHab operations: Danielle has taken charge of the GreenHab.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.